OpenStudy (anonymous):

Given f(x) = 4 x^2 + 9 x , determine the points where (a) the curve intercepts the x- axis; (b) the tangent line to the curve is horizontal. (a) The points where the curve intercepts the x- axis are . (The answers must be a POINTS on the curve; use parentheses for the points and commas to separate the points in the answer space.) (b) The point where the tangent line to the curve is horizontal is . (The answer must be a POINT on the curve; separate the values with a comma and use parantheses in the answer space.)

OpenStudy (anonymous):

a) This is a standard precalculus find the roots problem. Let f(x)=0 and solve. b) Same as the problem we did a minute ago.

OpenStudy (anonymous):

Make sure to show your answers in (x.y) form.

OpenStudy (anonymous):

ok. it confused me because its worded differently

OpenStudy (anonymous):

(1.125,15.1875) this is what i got, but its incorrect for the curve intercepts the x- axis;

OpenStudy (anonymous):

this is not correct :(

OpenStudy (anonymous):

hmm... did i did any calculation error?. the concept is there. weird..

OpenStudy (anonymous):

wait, i mean(0,0) and ( -9/4,0)

OpenStudy (anonymous):

(0,0),(-9/4,0)

OpenStudy (anonymous):

intercept x-axis, f(x)=0 0=4x^2 + 9x x=0, -9/4 when x=0, f(0)= 0 when x=-9/4, f(-9/4) =0 (0,0) and ( -9/4,0) dy/dx= 8x+9 horizoltal, dy/dx=0 8x+9=0 x=-9/8. when x=-9/8, f(-9/8)= -81/16 (-9/8, -81/16)

OpenStudy (anonymous):

yes thats it, thanks a million

OpenStudy (anonymous):

Remember, for a quadratic, you can find the vertex using the algebraic formula (-b/2a, f(-b/2a)). It won't win you many points as a calculus method, but is an excellent check step. Again, it only works on quadratics.

OpenStudy (anonymous):

thanks. :) everyone